# DIPLODOCUS I: Framework for the evaluation of relativistic transport equations with continuous forcing and discrete particle interactions

### v2

@@ II.3 Adding Collisions
III Coordinate-Dependent Transport Equation
III.1 Phase Flow with Coordinates
Though Eqs.˜9 and 10, are mathematically elegant in their coordinate-free form, some definition of coordinates (basis vectors) are needed to give them utility in modelling astrophysical sources.
For this purpose, two sets of basis vectors are used. The nature of spacetime is assumed to be a 4D Riemannian manifold $\mathcal{N}$, with signature (+++)(-+++), and a coordinate basis $\bm{e}_{\alpha}=$ (Greek scripts) with coordinates $x^{\alpha}=\{x^{0},...,x^{3}\}$ [^4] The choice of a precise set of coordinates, e.g. Cartesian, Cylindrical, Spherical, Schwarzchild, Boyer-Lindquist, etc. remains arbitrary. As such this model can be applied to a wide range of spacetimes.. For momentum space, which exists in the tangent space $T_{\mathcal{N}}$ to $\mathcal{N}$, a local-orthonormal basis (tetrad) $\bm{e}_{a}$ (Latin scripts), with components $p^{a}=\{p^{0},...,p^{3}\}$ is used. The worldline of a particle through the 8-dimensional phase space is then given by
@@ III.1 Phase Flow with Coordinates
$\begin{split}\mathcal{M}=&\{(\bm{x},\bm{p}):\bm{x}\in\mathcal{N},\bm{p}\in\bm{T}_{\mathcal{N}}\\ &\quad,\bm{p}^{2}\leq 0,\bm{p}\text{ future directed}\}.\end{split}$ (12)
The phase flow in $\mathcal{M}$ is generated by the Lie derivative with respect to $\bm{v}$, where, using Eq.˜11
@@
The total volume form is then given by the wedge product of Eqs.˜20 and 22:
@@
From Eq.˜23, the hypersurface element $\bm{\omega}$ orthogonal to the phase flow $\bm{v}$ (Eq.˜19) in $\mathcal{M}_{m}$ can also be evaluated:
@@
As discussed in Section˜II.2, for a system to be obey Liouville’s theorem $\bm{\mathrm{d}}\bm{\omega}\overset{!}{=}0$, by consideration of Eq.˜24 this condition is equivalent to $\frac{\partial F^{a}}{\partial p^{a}}\overset{!}{=}0$. As an example where this condition is met, the Lorentz force on a particle of charge $q$ and mass $m$ is given by $F^{a}=\frac{q}{m}F^{ab}p_{b}$, where $F^{ab}$ is the electromagnetic field tensor, therefore $\frac{\partial F^{a}}{\partial p^{a}}=F^{a}_{~a}=0$ as $F^{ab}$ is antisymmetric. However, a counter example is the radiation reaction force [^7] The radiation reaction force is also known as the Abraham-Lorentz-Dirac force [80, 81, 82] [see also 83, 84, for more modern descriptions]. experienced by a charged particle being accelerated by an electromagnetic field where $F^{a}\propto\left(\eta^{ab}+\frac{p^{a}p^{b}}{m^{2}c^{2}}\right)\left(\frac{q}{m}p^{d}p^{c}\partial_{c}F_{bd}+\frac{q^{2}}{m^{2}}qF_{bc}F^{cd}p_{d}\right)$, for such a force it can be shown that $\frac{\partial F^{a}}{\partial p^{a}}\neq 0$, and is therefore a non-conservative force, evident by the fact that it results in the charged particle losing energy over time.
As such, non-conservative forces, such as radiation reaction, are a necessary inclusion of any self-consistent astrophysical model, hence $\mathbf{d}\bm{\omega}\neq 0$ in general and the correct forms for the transport equation for the particle distribution function are Eqs.˜9 and 10.
III.2 Stationary Observers
@@
The hypersurface element $\bm{\omega}$, given by Eq.˜24, may be decomposed into hypersurface elements of simultaneous and non-simultaneous events as measured by a stationary observer:
@@
The first square bracketed term in Eq.˜25 is $\bm{\omega}_{\Sigma}$, referring to the element of $\Sigma$ hypersurfaces, while the second square bracket is the element of timelike hypersurfaces $\Lambda$ of non-simultaneous events, $\bm{\omega}_{\Lambda}$ (see Fig.˜5).
@@
With a set of observers, the total number of particles in the system can be counted in a well defined way. Consider the insight from Section˜II.2, that the $\int_{\partial Q}f\bm{\omega}$ is the number of worldlines passing through some hypersurface defined by the boundary $\partial Q$. Therefore, taking $\partial Q$ to be the open hypersurface $\Sigma$, $\int_{\Sigma}f\bm{\omega}=N$ the number of particles measured simultaneously passing through that surface, i.e. the total number of particles within the manifold at a given time $\tau$ according to that local observer. If $\Sigma=X\times P$ is decomposed into the product of a spatial submanifold $X$ and the momentum-space submanifold $P$ this integral can be expanded as
$N=\int_{\Sigma}f\bm{\omega}_{\Sigma}=\int_{X}-n_{a}n^{\alpha}\frac{\chi_{\alpha\beta\gamma\delta(\bm{x})}}{3!}\bm{\mathrm{d}}x^{\beta\gamma\delta}\int_{P}fp^{a}\bm{\pi}.$ (26)
The latter part of Eq.˜26 can be identified as the first moment of the distribution function i.e. locally measured 4-flow vector:
@@
An alternative to the set of Eulerian observers is a set of observers co-moving with some bulk flow. However, for co-moving observers, their local tetrad $\bm{e}_{a}$ is a function of the bulk flow, and therefore variable, unlike a Eulerian observer whose tetrad is fixed. This results in the hypersurface elements Eq.˜24 also becoming a function of the flow and therefore so too would the numerical fluxes (see Section˜IV.2) used to evaluate transport of the distribution function, thereby requiring unnecessary recalculation when the bulk flow changes.
III.3 The Collision Integral
The collision integral $\bm{C}(\bm{x},\bm{p})$ is a measure of the number of worldlines that are terminated and started within some volume of phase space. It is useful to relate this to the transition probability $w_{\mathfrak{if}}$ for an initial state $\mathfrak{i}$ of particles to transition to a final state $\mathfrak{f}$. Modifying the results of [85], the differential transition probability may be written:
@@ III.3 The Collision Integral
Given a distribution of particles $f(\bm{x},\bm{p})$, the number of particle worldlines crossing a $\Sigma$ hypersurface is given by Eq.˜26. Assuming a well defined basis transformation such $n_{a}=(-1,0,0,0)$, this may be integrated over a test region in the spatial sub-manifold $X$, giving
@@
where $V$ is some spatial volume. As such the expression $p^{0}V\bm{\pi}$ can be thought of as the momentum-space number density of particles with a unity distribution function, i.e. 1 particle per unit hypersurface element. The differential transition probability, Eq.˜30, may then be expanded to include the incident and final particle’s momentum states:
@@
Figure 6: The continuous distribution function $f$ for each particle is discretised via Eq.˜37 to form a range of plateaux over phase space (here taken to be two dimensional with coordinates $x$ and $y$) indexed by the values $f_{ij}$, hence the name Distribution-In-Plateaux.
@@
With the volume element for phase space given by Eq.˜23, it can be recognised that the quantity $\bm{N}_{\mathfrak{if}}$ is proportional to $\bm{\Omega}$, and therefore related to the collision integral $\bm{C}(\bm{x},\bm{p})$. To understand this relation, consider a particle of type-$\mathfrak{a}$. The number of its worldlines terminated by collisions where it is a member of the incident state $\mathfrak{i}$ is given by $\sum_{\mathfrak{a}\in\mathfrak{i}}\bm{N}_{\mathfrak{if}}$, and similarly the number of its worldlines that are started is $\sum_{\mathfrak{a}\in\mathfrak{f}}\bm{N}_{\mathfrak{if}}$. The nature of the collision integral for a single particle of type-$\mathfrak{a}$ is therefore
@@
Specific examples of the collision integral for binary (scattering) collisions i.e. $\mathfrak{ab}\rightarrow\mathfrak{cd}$ and emission (decay) processes i.e. $\mathfrak{a}\rightarrow\mathfrak{bc}$ are given in Appendix˜A.
Eq.˜36 contains integrals over the momentum states of all the initial and final state particles, therefore it is of high dimensionality and further complicated by potentially non-linear dependence on the particle distribution functions. As such, it is generally impractical to use in this form unless some methodology for simplifying the integral is applied, or some assumptions are made for the symmetries of $f$ in momentum space - traditionally, isotropy.
IV Distribution-In-Plateaux
The integrated form of the Boltzmann equations (Eq.˜10), is mathematically elegant but in its present form is difficult to evaluate. The difficulty arises from the integrals over the entire volume of phase space. This can be simplified by splitting the integration domain $Q$ into a set of continuous sub-domains with common boundaries. Therefore the left hand side of Eq.˜10 describes the flux of particle worldlines through the boundaries of each sub-domain, while the right hand side counts collisions within each sub-domain. As all boundaries are shared between sub-domains, the number of particle worldlines is conserved over the whole domain.
To aid in simplifying the equations in this section, the set of spacetime coordinates will be denoted by $x^{\alpha}=\{x^{0},x^{1},x^{2},x^{3}\}=\{t,x,y,z\}$, making the time coordinate $t$ explicit, and spatial coordinates $\{x,y,z\}$ refer, abstractly, to any set of three chosen spatial coordinates not just the standard Cartesian set i.e. $\{x,y,z\}$ (see Appendix˜B for examples). For momentum-space, coordinates are concretely defined to be $p^{i}=\{p^{1},p^{2},p^{3}\}=\{p,u,\phi\}$ the set of modified spherical coordinates ($u=\cos\theta$), as this is the set used for the numerical evaluation of collision integrals in Paper II.
With these coordinate sets, the procedure for splitting the integration domain $Q$ into a continuous set of sub-domains can formally be achieved by taking the particle distribution function to be constant within each sub-domain of phase space. The total distribution is formed by the sum of these constant regions [^8] It has been noted to the authors that this method of discretisation for particle transport may be mapped to a zeroth-order discontinuous Galerkin method, which was developed independently of this work, and further, has been shown to have good conservative properties for number density and energy (see, for example, Refs. [86, 87, 41]).:
@@ IV Distribution-In-Plateaux
where $H$ is a boxcar function and $\alpha,\beta,\gamma,\delta,i,j$ and $k$ now refer to a discrete set of phase-space coordinates. As an example, for the $t$ coordinate the function $H$ is defined as
@@
with $\Theta$ being the Heaviside step function and $h_{\alpha}^{\pm}$ being the boxcar functions values at the bounds of the domain $t_{\alpha}\leq t\leq t_{\alpha+1}$ [^9] By allowing the definition of $\Theta(0)$ to be a variable between 0 and 1, usually taken to be 0,1 or $1/2$, modification of the numerical scheme may be made, as described in Paper II, while maintaining conservative particle transport..
Through the action of Eq.˜37, the continuos distribution function is replaced by a series of plateaux, each with a “height” of $f_{\alpha\beta\gamma\delta ijk}$ over a discrete sub-domain in phase space (see Fig.˜6), and hence termed “Distribution-In-Plateaux” (DIP). This nomenclature has similarities with that of the “Particle-In-Cell” (PIC), whereby the distribution function is considered as a sum over Dirac delta functions.
The factor of $p^{2}\Delta p_{i}\Delta u_{j}\Delta\phi_{k}$ in Eq.˜37 is not strictly required, rather it has been introduced for convenience such that according to Eq.˜28 the scalar number density $n$ of particles measured by the stationary observer, within a spatial sub-domain $X_{\beta\gamma\delta}$ and between times $t_{\alpha}$ and $t_{\alpha+1}$, is given by
@@
Figure 7: Simplified schematic of how DIP applies to phase space transport (here phase space has been reduced to two dimensions). The action of Distribution-In-Plateaux Eq.˜37, is to discretise phase space into sub-domain in which the distribution function $f$ takes a single value ($f_{ij}$ here). The left hand side of the transport equation Eq.˜10 $\int_{\partial Q}f\bm{\omega}$ is then an integral over the boundaries of these sub-domains, which generates the fluxes between sub-domains ($\mathcal{A}$ and $\mathcal{B}$ here). For this two dimensional example, the the transport equation for $f_{22}$ takes the form $\int_{\partial Q_{22}}f\bm{\omega}=\left(h^{-}_{32}f_{32}+h^{+}_{22}f_{22}\right)\mathcal{A}^{+}+\left(h^{-}_{22}f_{22}+h^{+}_{12}f_{12}\right)\mathcal{A}^{-}+\left(h^{-}_{23}f_{23}+h^{+}_{22}f_{22}\right)\mathcal{B}^{+}+\left(h^{-}_{22}f_{22}+h^{+}_{21}f_{21}\right)\mathcal{B}^{-}$, where the $h$ terms (Eq.˜38) dictate the value of the distribution function on sub-domain boundaries.
IV.1 DIP Applied to Collision Integrals
Before describing how the DIP simplifies the larger problem of phase space transport. Consider how it simplifies the problem of evaluating collision integrals. For a binary collisions $\mathfrak{ab\rightarrow cd}$, the gain of particles of type-$\mathfrak{c}$ is given by the collision integral (Eq.˜36):
@@ IV.1 DIP Applied to Collision Integrals
where $G_{12\rightarrow 34}$ is the “gain term” given by Eq.˜51. The integration domain for this term is the phase space of particle $\mathfrak{c}$ and the momentum space of particles $\mathfrak{a}$ and $\mathfrak{b}$. By deploying DIP, the value of the distribution functions of particles $\mathfrak{a},\mathfrak{b}$, and $\mathfrak{c}$ become constant within a sub-domains of phase space and, therefore, can be taken outside integrals over those sub-domains. Consider a momentum-space sub-domain for particle $\mathfrak{c}$ to be $P_{ijk}=[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$, with DIP deployed, the collision integral (Eq.˜40) over this sub-domain becomes:
@@
where the gain term has become a 9-dimensional “gain array” $G_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d},ijklmnopq}$ given by Eq.˜53, and $\mathcal{V}_{\alpha\beta\gamma\delta}$ is the spacetime volume element given by Eq.˜71. The elements of the gain array are integrals over different momentum-space sub-domains of the incoming and outgoing particles, which, being independent of the distribution of those particles, are pre-computable by numerical means (see Paper II), and are proportional to the rate of a collision. Further, elements of the gain array are independent of spatial coordinates, due to the use of a local-orthonormal basis for momentum space, and as such are valid at all points in spacetime, independent of the spacetime’s nature. This process of discretisation and integration can be applied to all collision terms equally, thereby all collisions are presented as pre-computed arrays within the DIPLODOCUS framework.
The application of DIP to collision integrals presents some immediate advantages of using DIP over PIC. As all collision array terms may be pre-evaluated and reused over multiple simulations of phase-space transport, there is a reduction in computational intensity compared to the calculation of collisions during transport - especially when the number of particles (collisions) is large. In addition, it provides uniform sampling of collisions over the entire phase-space domain - independent of the number of particles in the system.
IV.2 DIP Applied to Transport

### v4

@@ 2.3 Adding Collisions
3 Coordinate-Dependent Transport Equation
3.1 Phase Flow with Coordinates
Though Eqs. (9) and (10), are mathematically elegant in their coordinate-free form, some definition of coordinates (basis vectors) are needed to give them utility in modelling astrophysical sources.
For this purpose, two sets of basis vectors are used. The nature of spacetime is assumed to be a 4D Riemannian manifold $\mathcal{N}$, with signature (+++)(-+++), and a coordinate basis $\bm{e}_{\alpha}=$ (Greek scripts) with coordinates $x^{\alpha}=\{x^{0},...,x^{3}\}$ [^4] The choice of a precise set of coordinates, e.g. Cartesian, cylindrical, spherical, Schwarzschild, Boyer-Lindquist, etc. remains arbitrary. As such this model can be applied to a wide range of spacetimes.. For momentum space, which exists in the tangent space $T_{\mathcal{N}}$ to $\mathcal{N}$, a local-orthonormal basis (tetrad) $\bm{e}_{a}$ (Latin scripts), with components $p^{a}=\{p^{0},...,p^{3}\}$ is used. The worldline of a particle through the 8-dimensional phase space is then given by
@@ 3.1 Phase Flow with Coordinates
$\begin{split}\mathcal{M}=&\{(\bm{x},\bm{p}):\bm{x}\in\mathcal{N},\bm{p}\in\bm{T}_{\mathcal{N}},\\ &\quad\bm{p}^{2}\leq 0,\bm{p}\text{ future directed}\}.\end{split}$ (12)
The phase flow in $\mathcal{M}$ is generated by the Lie derivative with respect to $\bm{v}$, where, using Eq. (11)
@@
The total volume form is then given by the wedge product of Eqs. (20) and (22):
@@
From Eq. (23), the hypersurface element $\bm{\omega}$ orthogonal to the phase flow $\bm{v}$ (Eq. (19)) in $\mathcal{M}_{m}$ can also be evaluated:
@@
As discussed in Section 2.2, for a system to obey Liouville’s theorem $\bm{\mathrm{d}}\bm{\omega}=0$ is required, by consideration of Eq. (24) this condition is equivalent to $\frac{\partial F^{a}}{\partial p^{a}}=0$. As an example where this condition is met, the Lorentz force on a particle of charge $q$ and mass $m$ is given by $F^{a}=\frac{q}{m}F^{ab}p_{b}$, where $F^{ab}$ is the electromagnetic field tensor, therefore $\frac{\partial F^{a}}{\partial p^{a}}=F^{a}_{~a}=0$ as $F^{ab}$ is antisymmetric. However, a counter example is the radiation reaction force [^7] The radiation reaction force is also known as the Abraham-Lorentz-Dirac force (Abraham, 1905; Lorentz, 1892; Dirac, 1938) (see also Gralla et al., 2009; Poisson et al., 2011, for more modern descriptions). experienced by a charged particle being accelerated by an electromagnetic field where $F^{a}\propto\left(\eta^{ab}+\frac{p^{a}p^{b}}{m^{2}c^{2}}\right)\left(\frac{q}{m}p^{d}p^{c}\partial_{c}F_{bd}+\frac{q^{2}}{m^{2}}qF_{bc}F^{cd}p_{d}\right)$, for such a force it can be shown that $\frac{\partial F^{a}}{\partial p^{a}}\neq 0$, and is therefore a non-conservative force, evident by the fact that it results in the charged particle losing energy over time.
As such non-conservative forces, e.g. radiation reaction, are a necessary inclusion of any self-consistent astrophysical model; hence $\mathbf{d}\bm{\omega}\neq 0$ in general and the correct forms for the transport equation for the particle distribution function are Eqs. (9) and (10).
3.2 Stationary Observers
@@
The hypersurface element $\bm{\omega}$, given by Eq. (24), may be decomposed into hypersurface elements of simultaneous and non-simultaneous events as measured by a stationary observer:
@@
The first square bracketed term in Eq. (25) is $\bm{\omega}_{\Sigma}$, referring to the element of $\Sigma$ hypersurfaces, while the second square bracket is the element of timelike hypersurfaces $\Lambda$ of non-simultaneous events, $\bm{\omega}_{\Lambda}$ (see Fig. 5).
@@
With a set of observers, the total number of particles in the system can be counted in a well defined way. Consider the insight from Section 2.2, that the $\int_{\partial Q}f\bm{\omega}$ is the number of worldlines passing through some hypersurface defined by the boundary $\partial Q$. Therefore, taking $\partial Q$ to be the open hypersurface $\Sigma$, $\int_{\Sigma}f\bm{\omega}=N$ the number of particles measured simultaneously passing through that surface, i.e. the total number of particles within the manifold at a given time $\tau$ according to that local observer. If $\Sigma=X\times P$ is decomposed into the product of a spatial submanifold $X$ and the momentum-space submanifold $P$ this integral can be expanded as:
$\begin{split}N&=\int_{\Sigma}f\bm{\omega}_{\Sigma}\\ &=\int_{X}-n_{a}n^{\alpha}\frac{\chi_{\alpha\beta\gamma\delta(\bm{x})}}{3!}\bm{\mathrm{d}}x^{\beta\gamma\delta}\int_{P}fp^{a}\bm{\pi}.\end{split}$ (26)
The latter part of Eq. (26) can be identified as the first moment of the distribution function i.e. locally measured 4-flow vector:
@@
An alternative to the set of Eulerian observers is a set of observers co-moving with some bulk flow. However, for co-moving observers, their local tetrad $\bm{e}_{a}$ is a function of the bulk flow, and therefore variable, unlike a Eulerian observer whose tetrad is fixed. This results in the hypersurface elements Eq. (24) also becoming a function of the flow and therefore so too would the numerical fluxes (see Section 4.2) used to evaluate transport of the distribution function, thereby requiring unnecessary recalculation when the bulk flow changes.
3.3 The Collision Integral
The collision integral $\bm{C}(\bm{x},\bm{p})$ is a measure of the number of worldlines that are terminated and started within some volume of phase space. It is useful to relate this to the transition probability $w_{\mathfrak{if}}$ for an initial state $\mathfrak{i}$ of particles to transition to a final state $\mathfrak{f}$. Modifying the results of Berestetskii et al. (1982), the differential transition probability may be written:
@@ 3.3 The Collision Integral
Given a distribution of particles $f(\bm{x},\bm{p})$, the number of particle worldlines crossing a $\Sigma$ hypersurface is given by Eq. (26). Assuming a well defined basis transformation $n_{a}=(-1,0,0,0)$ and Eq. (26) may be integrated over a test region in the spatial sub-manifold $X$, giving
@@
where $V$ is some spatial volume. As such the expression $p^{0}V\bm{\pi}$ can be thought of as the momentum-space number density of particles with a unity distribution function, i.e. 1 particle per unit hypersurface element. The differential transition probability, Eq. (30), may then be expanded to include the incident and final particle’s momentum states:
@@
With the volume element for phase space given by Eq. (23), it can be recognised that the quantity $\bm{N}_{\mathfrak{if}}$ is proportional to $\bm{\Omega}$, and therefore related to the collision integral $\bm{C}(\bm{x},\bm{p})$. To understand this relation, consider a particle of type-$\mathfrak{a}$. The number of its worldlines terminated by collisions where it is a member of the incident state $\mathfrak{i}$ is given by $\sum_{\mathfrak{a}\in\mathfrak{i}}\bm{N}_{\mathfrak{if}}$, and similarly the number of its worldlines that are started is $\sum_{\mathfrak{a}\in\mathfrak{f}}\bm{N}_{\mathfrak{if}}$.
The nature of the collision integral for a single particle of type-$\mathfrak{a}$ is therefore
@@
Specific examples of the collision integral for binary (scattering) collisions i.e. $\mathfrak{ab}\rightarrow\mathfrak{cd}$ and emission (decay) processes i.e. $\mathfrak{a}\rightarrow\mathfrak{bc}$ are given in Appendix A.
Eq. (36) contains integrals over the momentum states of all the initial and final state particles; it is therefore of high dimensionality and further complicated by potentially non-linear dependence on the particle distribution functions. As such, it is generally impractical to use in this form unless some methodology for simplifying the integral is applied, or some assumptions are made for the symmetries of $f$ in momentum space—traditionally, isotropy.
4 Distribution-In-Plateaux
Figure 6: The continuous distribution function $f$ for each particle is discretised via Eq. (37) to form a range of plateaux over phase space (here taken to be two dimensional with coordinates $x$ and $y$) indexed by the values $f_{ij}$, hence the name Distribution-In-Plateaux.
The integrated form of the Boltzmann equations (Eq. (10)) is mathematically elegant but in its present form is difficult to evaluate. The difficulty arises from the integrals being evaluated over the entire volume of phase space. This can be simplified by splitting the integration domain $Q$ into a set of continuous sub-domains with common boundaries. Therefore the left hand side of Eq. (10) describes the flux of particle worldlines through the boundaries of each sub-domain, while the right hand side counts collisions within each sub-domain. As all boundaries are shared between sub-domains, the number of particle worldlines is conserved over the whole domain.
[…]
With these coordinate sets, the procedure for splitting the integration domain $Q$ into a continuous set of sub-domains can formally be achieved by taking the particle distribution function to be constant within each sub-domain of phase space. The total distribution is formed by the sum of these constant regions:
@@ 4 Distribution-In-Plateaux
where indices $\alpha,\beta,\gamma,\delta,i,j,$ and $k$ now refer to a discrete set of phase-space coordinates and $H$ is a boxcar function. As an example, for the $t$ coordinate the function $H$ is defined as
@@
[…]
Through the action of Eq. (37), the continuous distribution function is replaced by a series of plateaux, each with a “height” of $f_{\alpha\beta\gamma\delta ijk}$ over a discrete sub-domain in phase space (see Fig. 6), and hence termed Distribution-In-Plateaux (DIP) [^9] It has been pointed out to the authors that this method of discretisation may be mapped to a zeroth-order discontinuous Galerkin method, which has been developed independently of this work, and further, has been shown to have good conservative properties for number density and energy (see, for example Cockburn and Shu, 2001; Gamba and Rjasanow, 2018; Gasparyan et al., 2022).. This nomenclature has similarities with that of the Particle-In-Cell (PIC), whereby the distribution function is considered as a sum over Dirac delta functions.
The factor of $p^{2}\Delta p_{i}\Delta u_{j}\Delta\phi_{k}$ in Eq. (37) is not strictly required, rather it has been introduced for convenience such that according to Eq. (28) the scalar number density $n$ of particles measured by the stationary observer, within a spatial sub-domain $X_{\beta\gamma\delta}$ and between times $t_{\alpha}$ and $t_{\alpha+1}$, is given by
@@
The application of DIP form of the distribution function to the equations of phase-space transport and collision integrals simplifies their evaluation in terms of expressions that can be implemented computationally. However, this comes at the cost of these expressions becoming rather long. These long expressions can be found in full in Appendices A and C, which will be referred to in the following sections.
4.1 DIP Applied to Collision Integrals
Before describing how DIP simplifies the larger problem of phase-space transport, it is useful to consider how it simplifies the problem of evaluating collision integrals. For a binary collision $\mathfrak{ab\rightarrow cd}$, the gain of particles of type-$\mathfrak{c}$ is given by the collision integral (Eq. (36)):
@@ 4.1 DIP Applied to Collision Integrals
where $G_{12\rightarrow 34}$ is the “gain term” given by Eq. (51) in Appendix A. The integration domain for this term is the phase space of particle $\mathfrak{c}$ and the momentum space of particles $\mathfrak{a}$ and $\mathfrak{b}$. By deploying DIP, the value of the distribution functions of particles $\mathfrak{a},\mathfrak{b}$, and $\mathfrak{c}$ become constant within a sub-domains of phase space and, therefore, can be taken outside integrals over those sub-domains. Consider a momentum-space sub-domain for particle $\mathfrak{c}$ to be $P_{ijk}=[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$, with DIP deployed, the collision integral (Eq. (40)) over this sub-domain becomes:
@@
[…]
The application of DIP to collision integrals presents some immediate advantages of using DIP over PIC. As all collision array terms may be pre-evaluated and reused over multiple simulations of phase-space transport, there is a reduction in computational intensity compared to the calculation of collisions during transport—especially when the number of particles (collisions) is large. In addition, it provides uniform sampling of collisions over the entire phase-space domain—independent of the number of particles in the system.
4.2 DIP Applied to Transport